A rectangle has a perimeter of 26 meters. Express the A (x) of the rectangle in terms of the length x of one of its sides.
Answer by fractalier(6550)

Well, perimeter can be found as P = 2L + 2W.
If the length is x, and the perimeter is 26, we have
26 = 2x + 2W
Now solve for W...
2W = 26 - 2x
W = 13 - x
Area is found by length times width...so
A(x) = LW = x(13-x)